The volume of gas at 65 kPa is 4.0 L. What is the volume of gas if the pressure increases to 105.0 kPa when amount and temperature are constant?


10. Which is larger? (circle/underline/highlight your answer)

2 bar compared to 1050 mmHg


11. Which is larger? (circle/underline/highlight your answer)

0.86 atm compared to 95.3 kPa


1
Expert's answer
2021-05-03T07:40:33-0400

1. Q.

The volume of gas at 65 kPa is 4.0 L. What is the volume of gas if the pressure increases to 105.0 kPa when amount and temperature are constant?




1. Answer = P1 V1 = P2 V2 .


65 × 4 = 105 × V2 .


V2 = 65 × 4 / 105 = 2.48 kpa .


P1 and V1 are initial pressure and Volume .


P2 and V2 are final pressure and Volume .



2. Q. Which is larger? (circle/underline/highlight your answer)

2 bar compared to 1050 mmHg .


Ans . 2 bar is higher .




Explanation =


1050 mm of Hg

atm = 1050 / 760

= 1.382 atm .

= 1.38 bar .




3. Q .

Which is larger? (circle/underline/highlight your answer)

0.86 atm compared to 95.3 kPa



3 Answer =

95.3 kPa have higher amount .


Explanation =

95.3 kPa =( 95.3 × 1000 ) Pa = 95300 Pa .



95300 Pa = 95300 × 1.01 × 105 atm .